# A Single Motor Nano Aerial Vehicle with Novel Peer-to-Peer Communication and Sensing Mechanism

### v2

@@ I Introduction
Figure 1: A photo of MP3. MP3’s size is $\varnothing 112\times 38$mm and weighs 20.0g. As MP3 flies, the whole robot rotates around the geometric axis indicated in the figure.
@@ II-B2 Optical Design
Figure 7: The receiver’s optical structure. (a) shows the structure of the left and right receivers, and (b) shows the structure of the middle receiver. There are two apertures in front of each photodiode to restrict the FOV to ±$\pm 10^{\circ}$ horizontal and ±$\pm 50^{\circ}$ vertical. The middle receiver’s aperture is tilted by $25^{\circ}$.
@@
Both the inner and outer apertures are important for creating a planar-like FOV. For the inner aperture, we want to increase its size to allow enough light to pass through. However, a finite-size inner aperture combined with a finite-size sensor creates a penumbra, where the light strength is reduced and varies depending on the direction. This introduces uncertainty about whether the signal will be received or not, which affects the accuracy of position sensing. Therefore, we also want to keep the inner aperture small to reduce this uncertainty. On the other hand, the outer aperture determines the FOV of the sensor, which influences how many transmitters could be seen at the same time and how long a transmitter will be in sight in a single rotation. These parameters are more important to communication and will be further explored in section II-B4. Eventually, we set the inner aperture size to $1.25\times 5$mm, horizontal FOV to ±$\pm 10^{\circ}$, vertical FOV to ±$\pm 50^{\circ}$, and the tilting angle of the aperture on the middle receiver to $25^{\circ}$.
@@ II-B3 Electrical Design
MP3 can only receive messages from others when its receivers are facing a transmitting MP3. MP3s typically spin at 25Hz and each of their receivers have a ±$\pm 10^{\circ}$ horizontal FOV, meaning that, in each revolution, a MP3 only has approximately 2-5ms to communicate with another MP3, and in the worst case scenario this gives a communication latency of 40ms. To facilitate communication with multiple MP3s using time division multiple access (TDMA), each MP3 could take no more than 500~1000us to transmit all necessary information to a receiver. If we want to communicate tens of bytes over this period, we need a system capable of communicating at around 1Mbps. This means that we could not use commercially available IR transceivers but had to implement our own.
Heavily constrained by space, weight, power, and limited available hardware solutions for optical communication, we decided to use a relatively simple communication method: transmitters can only transmit information by turning on and off LEDs, and the receivers will pass the analog received signal through a comparator and convert it back to a binary signal before being processed by the on-board microcontroller. With these constraints in mind, our transmitter and receiver circuit are designed to facilitate high frequency (-3dB analog bandwidth of 10.5M) modulation and reliable reception of IR signals in multi-point to multi-point communication scheme over $5-50$cm free space.
@@
It is worth noting that while the system is capable of communicating over 2m if we remove or enlarge the apertures, we want to limit the communication range to prevent congestion issues in communication. We want drones to be as close to each other as possible to create a dense swarm, so we expect robots to be around 30cm away from each other, and that motivates us to artificially limit the communication range to around 1.5×1.5$\times$1.5 × the average expected distance, or around 50cm.
@@ II-B4 Communication Protocol Design
Furthermore, a message interval of 50~100us means the resolution for timing measurements $t_{\text{1L}},t_{\text{1M}},t_{\text{1R}}$ is also 50~100us, which corresponds to an angle resolution of 0.5~$1^{\circ}$. Such angle resolution is just enough to obtain useful distance measurements (<2cm RMS error when the distance between MP3s is 30cm). If we increase the FOV of the receivers or reduce the time between transmissions, multiplexing ability will be reduced; and if we decrease the FOV or increase the time between transmissions, MP3s will not be able to transmit as much information in a single rotation. Furthermore, if we increase the time between transmissions, we will also lose position sensing accuracy.

### v3

@@ I Introduction
Figure 1: A photo of MP3. MP3’s size is $\varnothing 112\times 38$$m$ and weighs 20.0g. As MP3 flies, the whole robot rotates around the geometric axis indicated in the figure.
@@ II-B2 Optical Design
Figure 7: The receiver’s optical structure. (a) shows the structure of the left and right receivers, and (b) shows the structure of the middle receiver. There are two apertures in front of each photodiode to restrict the FOV to $\pm 10^{\circ}$ horizontal and $\pm 50^{\circ}$ vertical. The middle receiver’s aperture is tilted by $25^{\circ}$.
@@
Both the inner and outer apertures are important for creating a planar-like FOV. For the inner aperture, we want to increase its size to allow enough light to pass through. However, a finite-size inner aperture combined with a finite-size sensor creates a penumbra, where the light strength is reduced and varies depending on the direction. This introduces uncertainty about whether the signal will be received or not, which affects the accuracy of position sensing. Therefore, we also want to keep the inner aperture small to reduce this uncertainty. On the other hand, the outer aperture determines the FOV of the sensor, which influences how many transmitters could be seen at the same time and how long a transmitter will be in sight in a single rotation. These parameters are more important to communication and will be further explored in section II-B4. Eventually, we set the inner aperture size to $1.25\times 5$mm, horizontal FOV to $\pm 10^{\circ}$, vertical FOV to $\pm 50^{\circ}$, and the tilting angle of the aperture on the middle receiver to $25^{\circ}$.
@@ II-B3 Electrical Design
MP3 can only receive messages from others when its receivers are facing a transmitting MP3. MP3s typically spin at 25Hz and each of their receivers have a $\pm 10^{\circ}$ horizontal FOV, meaning that, in each revolution, a MP3 only has approximately 2-5ms to communicate with another MP3, and in the worst case scenario this gives a communication latency of 40ms. To facilitate communication with multiple MP3s using time division multiple access (TDMA), each MP3 could take no more than 500~1000us to transmit all necessary information to a receiver. If we want to communicate tens of bytes over this period, we need a system capable of communicating at around 1Mbps. This means that we could not use commercially available IR transceivers but had to implement our own.
Heavily constrained by space, weight, power, and limited available hardware solutions for optical communication, we decided to use a relatively simple communication method: transmitters can only transmit information by turning on and off LEDs, and the receivers will pass the analog received signal through a comparator and convert it back to a binary signal before being processed by the on-board microcontroller. With these constraints in mind, our transmitter and receiver circuit are designed to facilitate high frequency (-3dB analog bandwidth of 10.5M) modulation and reliable reception of IR signals in multi-point to multi-point communication scheme over 5505-50cm free space.
@@
It is worth noting that while the system is capable of communicating over 2m if we remove or enlarge the apertures, we want to limit the communication range to prevent congestion issues in communication. We want drones to be as close to each other as possible to create a dense swarm, so we expect robots to be around 30cm away from each other, and that motivates us to artificially limit the communication range to around $1.5\times$ the average expected distance, or around 50cm.
@@ II-B4 Communication Protocol Design
[…]